OpenStudy (anonymous):

Simplify each radical expression. Sqrt 45n^5. Pls help thnx..

OpenStudy (anonymous):

\[\sqrt{45n^5} = \sqrt{5*9 n^5} = \sqrt{5*9n^4*n}\] Then you get the 9n^4 out of the square root

OpenStudy (anonymous):

And then what next?

OpenStudy (anonymous):

What did you get by taking the 9n^4 out of the square root?

OpenStudy (anonymous):

I'm not getting it if u could explain pls thnx

OpenStudy (anonymous):

\[\sqrt{5*9n^4*n} = \sqrt{5n*9n^4} = \sqrt{5n} \sqrt{9n^4} = \sqrt{5n} \sqrt{(3n^2)^2}\] Then you simplify the second square root. Do you get it now?